OpenStudy (anonymous):

What is the product of 6the square root of 7. • 2the square root of 21.? Simplify if possible. 8the square root of 21. 12the square root of 3. 12the square root of 21. 84the square root of 3.

OpenStudy (anonymous):

@Bubblez<3

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ 6\sqrt{7}\times2\sqrt{21}= 6\times\sqrt{7}\times2\times \sqrt{21}= 6\times2\times\sqrt{7} \times\sqrt{21}=6\times2\times\sqrt{7\times21}=?}\)

OpenStudy (solomonzelman):

tell me if you need more help

OpenStudy (solomonzelman):

I was just playing with the order

OpenStudy (anonymous):

I do need help. lol

OpenStudy (anonymous):

@SolomonZelman